Peppermint Tea = 8.20/lb
Licorice Tea = 6.40/lb
How many pounds of each will result in 30 pounds of tea at 7.60/lb?
let lbs = x

Answer by mananth(16946)

price ---------------- quantity
Peppermint Tea 8.2 ---------------- x lb
Licorice tea 6.4 ------ 30 - x lb
Mixture 7.60 ---------------- 30
30
8.2 x + 6.4 ( 30 - x ) = 7.6 * 30
820 x + 640 ( 30 - x ) = 22800
820 x + 19200 - 640 x = 22800
820 x - 640 x = 22800 - -19200
180 x = 3600
/ 180
x = 20 lb 8.20% Peppermint Tea
10 lb 6.40% Licorice tea

m.ananth@hotmail.ca